### v6

# Random Schreier graphs as expanders

Abstract
Expander graphs, due to their good mixing properties, are useful in many algorithms and combinatorial constructions. One can produce an expander graph with high probability by taking a random graph. For example, for the case of bipartite graphs of degree $d$ and $n$ vertices in each part we may take independently $d$ permutations of an $n$-element set and use them for edges.
This construction is much simpler than all known explicit constructions of expanders and gives graphs with good mixing properties (small second largest eigenvalue) with high probability. However, from the practical viewpoint, it uses too many random bits, so it is difficult to generate and store these bits for reasonably large graphs. The natural idea is to replace the group of all permutations by its small subgroup. Let $n$ be $q^{k}-1$ for some $k$ and some prime $q$. Then we may interpret vertices as non-zero $k$-dimensional vector over the field $\mathbb{F}_{q}$, and take random linear permutations, i.e., random elements of $GL_{k}(\mathbb{F}_{q})$. In this way the number of random bits used will be polynomial in $k$ (i.e., the logarithm of the graph size, instead of the graph size itself) and the degree.
In this paper we provide some experimental data that show that indeed this replacement does not change much the mixing properties (the second eigenvalue) of the random graph that we obtain. These data are provided for several types of graphs (undirected regular and biregular bipartite graphs). We also prove some upper bounds for the second eigenvalue (though it is quite weak compared with the experimental results).
Finally, we discuss the possibility to decrease the number of random bits further by using Toeplitz matrices; our experiments show that this change makes the mixing properties of graphs only marginally worse, while the number of random bits decreases significantly.
Source Code:
1 Introduction
Consider a regular undirected graph with $n$ vertices and degree $d$, and a random walk: at every step, being at some vertex, we choose uniformly a random neighbor and go there. Then, after some fixed number of steps, we get a probability distribution on vertices. If this distribution converges fast to the uniform distribution as the number of steps increases, we say that the graph has good mixing properties: we will forget soon where we have started.
Adding one step to our walk means multiplying the probability distribution by the (normalized) adjacency matrix of the graph. The largest eigenvalue of this matrix corresponds to the stationary state (the uniform distribution), so the convergence speed is determined by the second largest eigenvalue of this matrix. Graphs having a good mixing property are called expanders. They are used in numerous applications in computer science and in coding theory. In these applications we usually need an instance of an expander graph.
How can we get a graph with good mixing properties, i.e., with small second eigenvalue (an expander)? The natural idea is to try a random graph. For example, we may consider $d/2$ random permutations of $n$ vertices (where $d$ is some even number) and use them as edges connecting the vertices before and after the permutation. Since we consider undirected graphs, we get a graph of degree $d$ (that might have parallel edges and loops). This contruction is called the permutation model. The largest eigenvalue equals $d$ (before the normalization that divides all elements by $d$), and all eigenvalues are real (since the matrix is symmetric). What can be said about the distribution of the second largest eigenvalue for a graph randomly sampled in this model?
This question was studied extensively (see [12] for a survey). In 1987 the bound $O(d^{3/4})$ (valid with high probability) was proven in [6]. Then it was improved by Friedman ([11], see also [4] for a shorter proof) who proved that (for every $\varepsilon>0$) the random construction gives the second eigenvalue at most $2\sqrt{d-1}+\varepsilon$ with high probability for all sufficiently large $n$ (number of vertices) and fixed $d$. This is rather close to the Alon–Boppana lower bound $2\sqrt{d-1}-\varepsilon$ that is valid (for every fixed $\varepsilon>0$) for all sufficiently large graphs ([20]), so the randomized construction is close to being optimal, at least asymptotically.
Observe also that the second largest eigenvalue for some particular graph can be computed efficiently, while a direct computation of the expansion ratio (as well as other combinatorial parameters characterizing the properties of expansion and connectivity) would require an exponential search over all subsets of the graph’s vertices. At the same time, we should notice that the spectral technique has its limitations, and computing the second largest eigenvalue cannot give better guarantee than an expansion ratio greater than half of the degree ([13]).
There are many different explicit constructions of expanders (see, e.g., [16, 9, 2, 19, 21, 18]); some of them also achieve a second eigenvalue close to the optimal. However, they are quite complicated, may have some restrictions on the degree, and the bounds for time complexity and eigenvalues can be of asymptotic nature. The advantage of the random graphs is the simplicity: we generate a graph using a simple randomized algorithm, check its second eigenvalue, and if it is small enough, we can use the graph when good expansion properties are needed. Note that the second eigenvalue needs to be computed only once (and this is still feasible for rather large graphs using modern algorithms and computers), and after that we may safely use the graph in our application.
It is important that the application does not need to recheck the eigenvalues. It still needs, however, to store the graph itself. For the random permutations model, it is a lot of information (each permutation of $n$ vertices requires about $n\log n$ bits). We may overcome this problem if we generate the random graph using much less random bits, store these bits in the application memory, and recompute the neighbors on the fly. But how can we decrease the number of random bits and still get good mixing properties?
In this paper we experiment with one approach of this type. Namely, we replace random permutations by linear random permutations. Let $\mathbb{F}_{q}$ be a finite field with $q$ elements. The $k$-dimensional vector space over this field contains $q^{k}$ elements ($k$-tuples of elements of $\mathbb{F}_{q}$). A $k\times k$ invertible matrix determines a permutation of these elements. The zero tuple is a fixed point, so we delete it and get a $q^{k}-1$-element set where the group $GL_{k}(\mathbb{F}_{q})$ of invertible $k\times k$ matrices over $\mathbb{F}_{q}$ acts transitively. The non-zero tuples will be the vertices of the graph, and $d/2$ invertible matrices give us $d/2$ permutations of vertices, so we may construct a regular graph with $q^{k}-1$ vertices and degree $d$.
The construction that we described is a special (and easy to implement) case of a Schreier graph. Let us recall the general definition.
Definition 1.1.
Let $H$ be a group acting transitively on a set $X$. We denote $h.x\in X$ the product of the action of $h\in H$ on $x\in X$. Let $S$ be a random multiset of elements of $H$ of $d$ elements. Then the Schreier (undirected) graph
$G=Sch(H\circlearrowleft X,S)$
is a $2d$-regular graph of size $|X|$ whose edges are $(x,s.x)$ for $x\in X$ and $s\in S$.
The graphs in our construction are Schreier graphs $Sch(H\circlearrowleft X,S)$ where $X=(\mathbb{F}_{q}^{k})^{*}$, $H=GL_{k}(\mathbb{F}_{q})$ acting on $X$ by the usual matrix-vector multiplication, and $S$ is a randomly chosen multiset of group elements.
This approach uses much less bits than the general permutation model: for the case $q=2$ we need about $k^{2}$ bits for each random matrix instead of $n\log n$ bits for a random permutation (and $n$ is about $2^{k}$). Increasing the size of the field, we get additional economy. The problem is that there are no theoretical guarantees for mixing properties: results proven for the random permutation model do not have yet their counterparts for random linear permutations (Though some weaker results exist; see [8], [22] or [10]). Note, however, that we can use this approach in practice safely, assuming we can check the eigenvalues before using the graph.
In this paper we provide some experimental evidence showing that this approach may work for several settings. In Section 2 we compare the empirical distributions for the second eigenvalue for random permutations and linear random permutations. We also compare different choices for the field size. In Section 3 we provide a similar comparison for bipartite graphs including the case where the sizes of the left and right parts differ significantly. To decrease the number of random bits further, we can use Toeplitz matrices; the experiments with them are reported in Section 4. Finally, in Section 5 we prove some upper bounds for the expected second largest eigenvalue of graphs considered in Section 2 and 3. Unfortunately, these bounds are quite complicated (the full proofs are given in the appendix) and quite weak; they are based on tools from [6] that give only $O(d^{3/4})$ bound. Still it can be considered as an additional motivation to consider only linear permutations instead of all permutations.
For reproducibility purposes, we provide the source code link in the front page.
2 Regular graphs
We consider a finite field $\mathbb{F}_{q}$ with $q$ elements, and the vector space of $k$-tuples $\mathbb{F}_{q}^{k}$ over $\mathbb{F}_{q}$. Then we choose randomly $d/2$ invertible matrices $T_{1},\ldots,T_{d/2}$, and use the corresponding permutations of the set of non-zero tuples to get an undirected graph of degree $d$. More precisely, a non-zero tuple $x\in\mathbb{F}_{q}^{k}$ has neighbors
$T_{1}x,\ldots,T_{d/2}x,T_{1}^{-1}x,\ldots,T_{d/2}^{-1}x.$
This is a regular graph with $q^{k}-1$ vertices and degree $d$ (that may contain loops and parallel edges), and we compute numerically the second eigenvalue of this graph.
This procedure is repeated many times, and then the empirical distribution of the second eigenvalues is shown (Figure 1). For comparison, we show also the distribution for the random permutation model, and for different field size.
Figure 1: To make the figure easier to read, all the eigenvalues measured have been divided by $2\sqrt{d-1}$. The two upper curves (quite close to each other) show the empirical distributions for random permutation model (dashed line) and random linear permutation model for graphs of degree $d=30$ ($15$ permutations) and size $116383=2^{14}-$ 1. The third curve (wider green one in the bottom part) shows the distribution for random linear permutations on $5$-tuples from a $7$-element field, so the graph has $168067^{5}-1=$ 16806 vertices (and the same degree $30$, obtained with $15$ permutations). For each distribution $5000$ experiments were made. The eigenvalues are grouped in $40$ bins
The invertible matrices were generated by choosing a random [^1] The random number generator used is Mersenne twister ([17]); we tested physical random generators without noticing any difference. matrix and then testing whether it is invertible. For Boolean matrices, the fraction of invertible ones is above $1/4$ (see, e.g., [1]), so we do not need too many trials; for bigger fields, the fraction is even bigger. Computing the determinant (checking the invertibility) is easy. For computing the eigenvalues we used the C++ library Spectra [^2] Spectra’s home page: https://spectralib.org/. that implements the Lanczos algorithm [7].
The choice of the unit on the horizontal axis ($2\sqrt{d-1}$ for non-normalized adjacency matrices) is motivated by the asymptotic lower bounds we mentioned. Graphs that have the second eigenvalue smaller than this unit are usually called Ramanujan graphs; we see that for our parameters both the random permutations and random linear permutations give Ramanujan graphs most of the time.
We also tested another set of parameters: we took $5$-tuples from a $7$-element field (so the resulting graph size is close to the previously considered one, $168067^{5}-1=$ 16806 instead of $163832^{14}-1=$ 16383, which makes the results roughly comparable). Note that this choice allows us to decrease the number of random bits used for the graph generation. We see that this is achieved for a price: the distribution is now less concentrated, though still very close to the Ramanujan threshold. (This is not a complete surprise that more random bits lead to better concentration.) One can also observe (we do not have any explanation for this paradoxical observation) that the probability to get a random graph that is on the left of, say, $0.998$, becomes bigger for a bigger field. So, if we try to get one graph with small eigenvalue by making many trials, the bigger field size may be preferable. One could try to go to the extreme in this direction and consider $1\times 1$ invertible matrices, i.e., consider multiplication by non-zero elements in some field $\mathbb{F}_{q}$ (the graph size is then $q-1$). This is, however, a bad idea since in this case the invertible matrices commute and therefore many paths in a random walk arrive at the same place just due to commutativity. Such a graph will have poor mixing properties.
Additionally, we can observe that variations on the degree of the graphs (for a given construction) does not seem to have much impact on the distribution of the second largest eigenvalues. We can observe this on Figure 2. Though the difference is hardly visible, we can see that the typical $10$-regular graph has slightly greater second largest eigenvalue compare to the unit we used ($2\sqrt{d-1}$). This is not surprising since low degree random graphs tend to be worse expanders. However, the eigenvalues are slightly more concentrated.
Figure 2: Second largest eigenvalue distribution of Schreier graphs of $GL_{14}(\mathbb{F}_{2})$ of degree $10,30,60$ and $120$.
3 Bipartite biregular graphs
We have discussed undirected regular graphs with good mixing properties, but in some cases one may wish to construct a bipartite graph that is biregular (degrees of the vertices are the same in each part, but may differ in left and right parts) and has good mixing properties (for the back and forth random walk).
Note first that every undirected regular graph can be converted to a bipartite biregular graph (with the same degree in both parts), just by doubling the set of vertices (each vertex now is present both in the left and right part) and interpreting the edges as edges between parts. The random walk in this graph is just the random walk in the original undirected graph (combined with alternating the parts), so the mixing properties of this bipartite graph are the same as for the initial undirected regular graph.
However, in the bipartite case we do not need to use permutation together with its inverse. The random permutation model thus can be modified: we consider $d$ random permutations instead; each one determines a one-to-one correspondence between parts and we draw edges that represent these $d$ correspondences. We use this construction in our experiments with bipartite graphs (the previous construction would have given the same result as for regular non bipartite graphs). This allows to have bipartite graphs of any (possible odd) degree.
Sometimes (e.g., in some constructions in coding theory) we need biregular bipartite graphs where left and right parts are of different size. A natural way to get such a graph is to start with a bipartite graph with the same part sizes, and then merge vertices in one of the parts. Merging $s$ vertices into one, we decrease the size and increase the degree by the factor of $s$. (We assume that number of vertices is a multiple of $s$.)
These constructions can be performed both with random permutation and random linear permutations. The same question arise: what can we say about mixing properties of graphs in these two settings? A traditional way to measure mixing property of a bipartite graph is to consider the (non-normalized) adjacency matrix for it (height and width are equal to the total number of vertices in both parts). The eigenvalues of this matrix are grouped into pairs, starting with $\sqrt{d_{L}d_{R}}$, where $d_{L}$ and $d_{R}$ are left and right degrees (the eigenvector is $\sqrt{d_{L}}$ in the left part and $\sqrt{d_{R}}$ in the right part) and $-\sqrt{d_{L}d_{R}}$ (to change the sign of the eigenvalue we change signs in all eigenvector coordinates in one of the parts). The next pair, $\lambda>0$ and $-\lambda<0$, determines the mixing properties, and we call $\lambda$ the second largest eigenvalue (ignoring the negative ones).
The role of the Ramanujan threshold $2\sqrt{d-1}$ (for regular undirected graphs of degree $d$) is now played by $\sqrt{d_{L}-1}+\sqrt{d_{R}-1}$: it has been proven that the second largest eigenvalue cannot be much smaller than this quantity [15] and with high probability is not much larger for random graphs [5]. We use $\sqrt{d_{L}-1}+\sqrt{d_{R}-1}$ as the unit on the horizontal axis for all our figures.
Figure 3: Here, n1 is the size of each partition. The dashed line (blue) shows the distribution of second eigenvalues for a bipartite graph constructed from $30$ random permutations of $16383$-element set (nonzero bit vectors of size $14$). The orange line (close to the first one) shows the same for $30$ linear permutations (over $\mathbb{F}_{2}$). The third line (green, in the bottom part) shows the distribution for $30$ permutations of a $16806$-element set randomly chosen among the $\mathbb{F}_{7}$-linear permutations of nonzero $5$-tuples with elements in $\mathbb{F}_{7}$.
Figure 3 shows the experiments with bipartite graphs with the same degree $30$ in both parts obtained using $30$ permutations of a set of non-zero elements of $\mathbb{F}_{2}^{14}$. The setting is different from the case of the undirected regular graph (Figure 1) where we used $15$ permutations and their inverses, but the results turn out to be quite similar (and the distribution for the case of $\mathbb{F}_{7}^{5}$ is also similar to Figure 1).
For the case of different left and right degrees the results of our numerical experiments are shown in Figure 4.
Figure 4: The dashed line (blue) shows the distribution of the second eigenvalues for $5000$ experiments when we take $10$ random permutations of a set with $163832^{14}-1=$ 16383 elements, construct a bipartite graph with degree $10$ in both sides and then merge triples of vertices on one side thus getting degree $30$ and size $16383/3=5461$. The solid (orange) line replaces random permutations by random linear (over $\mathbb{F}_{2}$) permutations of non-zero vectors in $\mathbb{F}_{2}^{14}$. Finally, the line with dots (green) shows the same distribution if we take $10$ linear permutations of non-zero vectors in $\mathbb{F}_{7}^{5}$.
Comparing this figure with the previous one, we observe some differences (that we cannot explain): the random permutations now give slightly (but visibly) better results compared to random linear permutations. On the other hand, changing the size of the field (and matrix dimension) now has much less effect. In fact, results may depend on which triples we are merging, for the linear case. The choice of the merged triplets of vertices could have some influence in the linear permutation case. This might be due to the additional structure induced by the the construction (compared to general permutations). However, this phenomenon remains to be explained.
4 Graphs from Toeplitz matrices
We now turn to modify our construction in order to reduce further the number of used random bits. Instead of taking random invertible matrices, we take Toeplitz matrices. Recall that a matrix $T=(t_{ij})$ is called a Toeplitz matrix if $t_{ij}$ depends only on the difference $j-i$, i.e., if
$T=\begin{pmatrix}a_{0}&a_{1}&a_{2}&\cdots&\cdots&\cdots&a_{k-1}\\ a_{-1}&a_{0}&a_{1}&\ddots&&&\vdots\\ a_{-2}&a_{-1}&a_{0}&\ddots&\ddots&&\vdots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ \vdots&&\ddots&\ddots&a_{0}&a_{1}&a_{2}\\ \vdots&&&\ddots&a_{-1}&a_{0}&a_{1}\\ a_{-k+1}&\cdots&\cdots&\cdots&a_{-2}&a_{-1}&a_{0}\\ \end{pmatrix}$
for some $a_{-k+1},\ldots,a_{0},\ldots,a_{k-1}$.
If we choose the elements of such a matrix randomly from a field of characteristics $q$, then we get an invertible matrix with probability $1-1/q$ (see [14]), so it is even easier to generate them. A Toeplitz matrix is defined by $12k-$ 1 elements of the field (instead of $k^{2}$ for a general matrix), so this is also an advantage. (One more technical advantage is that we can multiply a vector by a Toeplitz matrix very fast, since it is essentially the convolution.)
Figure 5: Empirical distributions of the second largest eigenvalue of 5000 graphs obtained from Toeplitz matrices (labeled “TP’) (for $\mathbb{F}_{2}^{14}$ and $\mathbb{F}_{7}^{5}$) compared to the corresponding empirical distributions for random linear permutations in the same vector spaces (labeled “GL”).
In Figure 5 we return to the case of undirected regular graphs, and compare random graphs obtained using $15$ random matrices (either random invertible matrices or random invertible Toeplitz matrices). Two upper curves (yellow and blue) use $\mathbb{F}_{2}^{14}$; one can see that Toeplitz matrices give bigger (and less concentrated) eigenvalues. The same can be seen from two bottom curves (red and green); we see also that the decreasing the dimension and increasing field size increase significantly the variance (but this is not the case for the mean in our example).
5 Computer-assisted theoretical bounds
The numerical experiments show that the constructions discussed in the previous sections allow to sample good spectral expanders of reasonably small size (e.g., with $n\sim 10^{4}$ vertices, when we can keep the entire graph in the memory of a computer). This might be enough for practical applications: we sample at random a multiset $S$ of matrices, produce the corresponding graph, and compute its eigenvalues. If we obtain a good spectral expander, we can fix the used random bits and plug this instance of a graph in our application.
However, in some applications we may need “strongly explicit” expanders with much larger number of vertices (e.g., with $n\sim 2^{200}$). In this case we cannot produce the entire matrix of the graph, although we still can keep in the memory an index of one vertex and we can compute efficiently the indices of its neighbours. In this case, we are not able to compute numerically the spectrum of the graph. So it would be helpful to have a theoretical result saying that most of (pseudo)random graphs (for some model of sampling) are good spectral expanders. Another reason why we are interested in theoretical bounds (even for graphs of small size) is less practical: we would like to explain the phenomena that we observe in numerical experiments.
In what follows we adapt to our setting the proof technique from [6] (we use the trace method and analyse the probability of getting a closed path during a random walk). We use it to prove bounds on the expected second largest eigenvalue of the adjacency matrix of a sampled graph. Our proofs involve heavy computations, so they are in some sense computer-assisted. All our theoretical results (except Theorem 3) only apply to $\mathbb{F}_{2}$ (though the experimental results show that the construction works well for bigger fields, see Figures 1 and 3).
5.1 Spectral bound for regular graphs
To simplify the notations, here $d$ will represent the number of random matrices used. Hence the degree of the graph will be $2d$.
Theorem 1.
Let $G$ be a Schreier graph of $GL_{k}(\mathbb{F}_{2})$ acting on $(\mathbb{F}_{2}^{k})^{*}$ with the random multiset $S$, $n=2^{k}-1$ (with $k\geq 2$), and degree $2d=2|S|$. Let $\mu_{2}$ be the second largest eigenvalue of its normalized adjacency matrix. Consider the following recurrent relation:
$X_{p}(c,\ell,d)=\left\{\begin{array}[]{ll}1\mbox{ if }c=0\mbox{ and }\ell=0;\\ 0\mbox{ if }p>\ell;\\ \sum_{i=c}^{\lfloor\frac{\ell}{p}\rfloor}\binom{d}{i}\bigg{(}\frac{2^{p}}{p!}\bigg{)}^{i}\frac{\ell!}{(\ell-ip)!}X_{p+1}(0,\ell-ip,d-i)\mbox{ otherwise.}\end{array}\right.$
We set $x_{3}(i)=X_{3}(1,2m-2i,d-i)$ and $x_{4}(i)=X_{4}(0,2m-2i,d-i)$. Then, for every integer $m$,
$\mathbf{E}(|\mu_{2}|)\leq\\ \bigg{(}\bigg{(}\frac{1}{2d}\bigg{)}^{2m}n\bigg{[}\sum_{i=1}^{m}\binom{d}{i}\frac{(2m)!}{(2m-2i)!}\bigg{[}(2^{i}-1)(x_{3}(i)+x_{4}(i))\frac{2}{n}\\ +\bigg{(}(x_{3}(i)+x_{4}(i))\frac{1}{n}+\frac{2^{i}}{(i+1)!}\bigg{(}x_{3}(i)\frac{5}{n}+x_{4}(i)\bigg{)}\bigg{)}\bigg{]}\\ +X_{1}(1,2m,d)\frac{1}{n}+x_{3}(0)\frac{5}{n}+x_{4}(0)\bigg{]}-1\bigg{)}^{\frac{1}{2m}}.$ (1)
The bound (1) looks messy, and the asymptotic analysis may be difficult. However, this formula becomes useful when we need to calculate (with the help of a computer) a bound for a specific graph of relatively small size (e.g., less than $2^{200}$ vertices).
To apply this theorem for given $k$ and $d$, we choose the value of $m$ such that (1) implies the strongest bound for $\mathbf{E}(|\mu_{2}|)$. We discovered that for several “reasonable” values of $k$ and $d$, with $2m$ chosen close to $k$, this formula implies a non trivial bound for $\mu_{2}$.
5.2 Spectral bound for bipartite graphs
Theorem 2.
Let $G$ be a Schreier bipartite graph of $GL_{k}(\mathbb{F}_{2})$ ($k\geq 2$) acting on $(\mathbb{F}_{2}^{k})^{*}$ with the random multiset $D$, $|D|=d$, and $n=|(\mathbb{F}_{2}^{k})^{*}|$. Let $\mu_{2}$ be the second largest eigenvalue of its normalized adjacency matrix. Consider the relation
$Y_{p}(c,\ell,d)=\left\{\begin{array}[]{ll}1\mbox{ if }c=0\mbox{ and }\ell=0;\\ 0\mbox{ if }p>\ell;\\ \sum_{i=c}^{\lfloor\frac{\ell}{p}\rfloor}\binom{d}{i}(p!)^{-i}\frac{\ell!}{(\ell-ip)!}Y_{p+1}(0,\ell-ip,d-i)\mbox{ otherwise.}\end{array}\right.$
We set $y_{3}(i)=Y_{3}(1,2m-2i,d-i)$, $y_{4}(i)=Y_{4}(0,2m-2i,d-i)$. Then, for every integer $m$,
$\mathbf{E}(|\mu_{2}|)\leq\bigg{(}\bigg{(}\frac{1}{d}\bigg{)}^{2m}n\bigg{[}\bigg{(}\frac{1}{d}\bigg{)}^{2m}\\ \sum_{i=1}^{m}\bigg{[}\binom{d}{i}\bigg{(}\frac{1}{2}\bigg{)}^{i}\frac{(2m!)}{(2m-2i)!}\frac{2^{i}}{(i+1)!}\bigg{(}y_{3}(i)\frac{5}{n}+y_{4}(i)\bigg{)}\bigg{]}\\ +Y_{1}(1,2m,d)\frac{1}{n}+Y_{2}(1,2m,d)\frac{2}{n}+y_{3}(0)\frac{5}{n}+y_{4}(0)\bigg{]}-1\bigg{)}^{\frac{1}{2m}}.$ (2)
Again, the expression in (2) is difficult to analyse, but it helps to compute specific (not asymptotic) bounds for some specific values of parameters.
Finally, in order to get theoretical guarantees on biregular graphs, we also prove this upper bound on the merging operation of bipartite regular graphs. Here we consider the non normalized eigenvalues.
Proposition 1.
Let $G$ be the bipartite $d_{1}$-regular graph and let $n$ be the size of each partition. We denote by $\alpha$ some upper bound on the second largest eigenvalue of $G$. Let $\gamma$ be an integer that divides $n$ and set $d_{2}=\gamma d_{1}$. Let $H$ be the bipartite $d_{1}d_{2}$-biregular graph obtained by merging every $\gamma$ vertices of one of the partitions of $G$ (in an arbitrary order). Let $\lambda_{2}$ be the second largest eigenvalue of $H$. Then,
$|\lambda_{2}|\leq\sqrt{d_{1}d_{2}}\alpha.$
5.3 A simple asymptotic bound
In this work we do not focus on asymptotic behaviour of large graphs. However, we observe that at least in some setting, the used technique easily gives some asymptotic bounds. This bound applies to the construction of a Schreier graph $GL_{k}(\mathbb{F}_{q})$ with any dimension $k$ (including $k=1$) and with any prime $q$, which is not the case of the previous ones (in Theorem 1 and Theorem 2 we use very substantially the fact that the group is not abelian, which is false in dimension one). The bound applies to graphs with a pretty large degree, $d=\Omega(\ln n)$ (see next section for an estimate of the constant hidden in the $\Omega$ notation).
Theorem 3.
Let $G=Sch(GL_{k}(\mathbb{F}_{q})\circlearrowleft(\mathbb{F}_{q}^{k})^{*},S)$ graph of size $n=q^{k}-1$ and whose degree is $2d=2|S|$. Let $\mu_{2}$ be the second largest eigenvalue of its normalised adjacency matrix. Then,
$\mathbf{E}(|\mu_{2}|)\leq e\sqrt{\frac{\ln n}{d}}$
where $\ln$ is the natural logarithm and $e$ its base.
5.4 Computer-assisted upper bounds for the second eigenvalue
Our experimental observations show that the second eigenvalue of a random Schreier graph from our construction is very close to the theoretical optimum (among all regular graphs). This phenomenon remains unexplained. What is even more frustrating, and we cannot guarantee that (pseudo) random graphs from similar families but with a large number of vertices (e.g., with $n\sim 2^{200}$) have good spectral properties. However, Theorem 1 and Theorem 2 imply some nontrivial bounds for the second eigenvalues of those graphs. In Table 1 we show several examples of such bounds and compare them with the Ramanujan’s threshold.
We observed that the bound from Theorem 1 becomes non-trivial (i.e., the bound for the normalised second eigenvalue $\mu_{2}$ is less than $1$) for $d=\Omega(k)$ ($d$ is the degree of the graph). The experiments suggest that for $d=2k$ and large number of vertices this theorem gives a bound that converges to a constant around $0.7$ (although we do not have a formal justification of this hypothesis). This bound gives more interesting implications for larger degree of the graph (compared with parameter $k$, e.g., for $d\geq k^{2}$). If we compare the bounds from Theorem 1 and 2 in terms of number of random matrices used, the first bound could seem slightly better than the second (though pretty similar). However, if we compare them in terms of degree, Theorem 2 gives a better bound, see Table 1. This is because increasing the number of matrices used improves the bound in both theorems and we need twice as much random matrices for bipartite graphs to get the same degree.
$(k,d)$ Th.1 ratio Th.2 ratio
$(14,30)$ 0.7607 2.1190 0.5787 1.6122
$(14,60)$ 0.5033 1.9657 0.3923 1.5325
$(14,120)$ 0.3389 1.8640 0.2687 1.4781
$(20,40)$ 0.7758 2.4847 0.5741 1.8387
$(25,100)$ 0.4969 2.4971 0.3718 1.8685
$(30,1000)$ 0.1435 2.2709 0.1128 1.7846
$(40,1000)$ 0.1618 2.5604 0.1251 1.9792
$(60,120)$ 0.7377 4.0577 0.5086 2.7977
$(200,400)$ 0.7016 7.0251 0.4592 4.5977
Table 1: This table shows the upper bound for the expected value of $\mu_{2}$ computed with Theorems 12 for a few examples of parameters values (the dimension $k$ and the degree $d$). Note that the regular non bipartite graphs are constructed with $d/2$ random matrices whereas the bipartite ones used $d$ random matrices. We display also the ratio between the Ramanujan’s threshold and the values given by the bounds ($\mu_{2}/\frac{2\sqrt{d-1}}{d}$). The first three lines correspondent to graphs that we could test in our experiments
6 Final comments
In this paper we study a pseudo-random construction of spectral expanders represented as Schreier graphs. The experimental results suggest that these graphs have nearly optimal value for the second largest eigenvalues (close to $2\sqrt{d-1}$ for regular graphs of degree $d$ and, respectively, to $\sqrt{d_{1}-1}+\sqrt{d_{2}-1}$ for bi-uniform bipartite graphs with degrees $d_{1}$ and $d_{2}$), and this is true not asymptotically but for graphs of reasonably small sizes relevant for practical applications.
A possible direction of the future research is a reduction of the number of random bits used to produce each graph. It would be interesting to reduce the number of these bits from $O(d\log^{2}n)$ to $O(d\log n)$ and improve computational complexity of the construction. Our experiments show that a promising direction is a construction based on Toeplitz matrices.
The “computer assisted theoretical bounds” that we provide are a small step to an explanation of the obtained experimental results. Instead of more traditional asymptotic bounds, we focused on theoretical bounds that can be calculated (possibly with the help of computer) for graphs with specific values of parameters, including the values of parameters that appear in our numerical experiments. In the appendix (A.4), we discuss limitations of the used method. It seems that a better theoretical explanation of the behaviour of random Schreier graphs of $GL_{k}(\mathbb{F}_{q})$ will require a subtler analysis.
We would also like to draw attention to the observation that useful theoretical bounds and estimates can be obtained with the help of “hideous” formulas combined with computer computations. This approach contradicts the spirit of classical mathematics, where we appreciate proofs that are elegant and meaningful for a human. However, these techniques can be useful to justify properties of objects (e.g., of pseudo-random graphs) with specific parameters, which may be necessary in practical applications.
Acknowledgements
The author thanks his PhD advisers, Andrei Romashchenko and Alexander Shen, for providing several ideas, useful remarks and help on this work.
Appendix A Proofs of the main results
A.1 Proofs of Theorems 1 and 3
Let $q$ be a prime, $k$ a natural integer and $G=Sch(GL_{k}(\mathbb{F}_{q})\circlearrowleft(\mathbb{F}_{q}^{k})^{*},S)$ be a Schreier graph with $S$ a random subset of $d$ elements of $GL_{k}(\mathbb{F}_{q})$ and their inverses. Hence, we obtain a $2d$ regular graph whose size is $n=q^{k}-1$. By mapping every element of $(\mathbb{F}_{q}^{k})^{*}$ to an element of $[\![1,\ q^{k}-1]\!]$ (e.g. with $f:(x_{1},\dots,x_{k})\mapsto\sum_{i=0}^{k-1}x_{i}q^{i}$), we can associate every pair of vertices with a coordinate of a matrix. This way, we can define $M$, the normalised adjacency matrix of $G$. Let
$1=|\mu_{1}|\geq|\mu_{2}|\geq...\geq|\mu_{n}|$
be its eigenvalues (which are real since the matrix is symmetric).
Consider a random walk of $2m$ steps starting at vertex $i$. Then the $(i,i)$ coordinate of $M^{2m}$ corresponds to the number of closed walks starting at vertex $i$ of size $2m$ divided by $(2d)^{2m}$, since $(2d)^{2m}$ is the number of paths of size $2m$ starting at $i$. Therefore, this is the probability (denoted $P_{ii}$) of returning to the vertex $i$ after $2m$ steps of the random walk. Since $Trace(M^{2m})$ is equal to the sum of all of these quantities and since the expected $(i,i)$ coordinate of $M^{2m}$ is the same for every $i$ we get
$\mathbf{E}(Trace(M^{2m}))=n\mathbf{E}(P_{11}).$
On the other hand, we have
$\sum_{i=1}^{n}\mu_{i}^{2m}=Trace(M^{2m}).$
Thus, since $\mu_{i}^{2m}\geq 0$ and $\mu_{1}=1$, we get $|\mu_{2}|\leq(Trace(M^{2m})-1)^{1/2m}$, which implies by Jensen’s inequality
$\mathbf{E}(|\mu_{2}|)\leq(n\mathbf{E}(P_{11})-1)^{1/2m}$ (3)
Given the starting vertex (say $v_{1}=f^{-1}(1)$), the random walk can be seen as the product of $2m$ elements of $S$ chosen uniformly and independently at random. We denote this product $\omega=w_{2m}.w_{2m-1}\dots w_{2}.w_{1}$ and each of its matrices represents the choice of a particular neighbour for all vertices. Hence, if the first vertex on the path is $v_{1}$, the second will be $v_{2}=s_{i_{1}}.v_{1}$ (where $s_{i_{1}}$ is the value of $w_{1}$), the third $v_{3}=s_{i_{2}}.v_{2}$ (where $s_{i_{2}}$ is the value of $w_{2}$) and so on; the last vertex of the path is then $\omega.v_{1}$.
We reuse here the main conceptual idea of the proof in [6]. The value of $\omega.v_{1}$ depends on two types of random choices: on the random choice of the word $\omega=w_{2m}.w_{2m-1}\dots w_{2}.w_{1}$ where each letter is chosen at random in $\{s_{1},s^{-1}_{1},\dots,s_{d},s_{d}^{-1}\}$, and the random choice of a matrix in $GL_{k}(\mathbb{F}_{q})$ for each $s_{i}$. These two choices are independent. We may sample at first the words $\omega$ and only then choose the matrices $s_{j}$. We prefer not to sample the entire value of each $s_{j}$ in “one shot” but reveal the values of these matrices (better to say, the values of the linear operators corresponding to these matrices) little by little, as it is needed. Thus, starting at vertex $v_{1}$, instead of choosing at random in $GL_{k}(\mathbb{F}_{q})$ the entire matrix $w_{1}=s_{i_{1}}$, we only determine the result of the product $v_{2}=s_{i_{1}}.v_{1}$. This choice does not determine completely the matrix $s_{i}$ but imposes a linear constraint on the matrix elements of $s_{i_{1}}$. The same letter $w_{1}$ may appear in the word $\omega$ several times. Each time the same letter $w_{1}$ appears in the word $\omega$ and, therefore, the matrix $s_{i_{1}}$ is encountered on the path, we must define the action of this matrix on some new vector $x$. We choose the result of $s_{i_{1}}.x$ by extending the partial definition of $s_{i_{1}}$, which means an extension of the linear constraints on $s_{i}$ fixed earlier. In a similar way, we define step by step the other matrices $s_{j}$ that are involved in $\omega$. We need to understand the distribution of the vector $v_{2m}=\omega.v_{1}$ that we obtain at the end of this procedure (and the probability of the event $v_{2m}=v_{1}$). In the next paragraphs we analyse this distribution.
Consider a matrix $s\in S$ that has been already encountered on the path refined by $\omega$, and we have already defined the action of $s$ on $t$ different vertices. Assume that we encounter the same matrix $s$ once again, and we must define the product $s.x$ for some one more vector $x\in(\mathbb{F}_{q}^{k})^{*}$). In the permutation model, as stated in [6] this would be a uniform distribution over the $n-t$ vertices that have not been earlier assigned to the partially defined permutation $s$. However, in our construction, even if $x$ is totally new to $s$, the result of $s.x$ may not be necessarily undetermined. Indeed, if $x$ is linearly dependent from the vectors that we have already met, we would have
$x=\sum_{i=1}^{t}\alpha_{i}x_{i},$
which is a sum of vectors whose result, when multiplied by $s$, is already known. Thus,
$s.x=\sum_{i=1}^{t}\alpha_{i}s.x_{i}$
would be completely determined by our previous random choices, and would not give any new information about $s$. Intuitively, we would say that the step that leads from $x$ to $s.x$ is not free. In order characterise formally what it means for a step to be free, we need to introduce the following set: let $s$ be a matrix of $S$, $\omega=w_{2m}\dots w_{1}$, $v_{1}$ the starting vertex and $v_{j+1}=w_{j}v_{j}$. Then we define
$\Sigma_{s}(i)=\operatorname{span}(\{v_{j}:j<i,w_{j}=s\}\cup\{v_{j+1}:j<i,w_{j}=s^{-1}\}).$
This is the set of vector on which the action of $s$ is determined at step $i$. The image set is thus
$s.\Sigma_{s}(i)=\operatorname{span}(\{v_{j+1}:j<i,w_{j}=s\}\cup\{v_{j}:j<i,w_{j}=s^{-1}\}).$
This leads to the analogous definition that is presented in [6].
Definition A.1 (free and forced step).
We consider the $i$-th step in the path. Let $s=w_{i}$. We say that step $i$ is forced when $v_{i}\in\Sigma_{s}(i)$. In the opposite case, we say that the step $i$ is free.
Alternatively, instead of saying that a step $i$ is free, we will say that the vector obtained after this step is free (namely the $i+1$-th vector, $w_{i}.v_{i}$). The following lemma justifies this terminology, and will be used systematically in the rest of the paper.
Lemma A.1.1.
Let $s=w_{i}$ for a step $i$ and $t$ be the dimension of $\Sigma_{s}(i)$. Then, if $v_{i}\notin\Sigma_{s},\ v_{i+1}=s.v_{i}$ can be chosen uniformly at random among the $q^{k}-q^{t}$ vectors that do not belong to $s.\Sigma_{s}(i)$.
Proof.
The choice of a non degenerate matrix $s$ of size $k\times k$ is the same as the choice of an bijective linear operator from $\mathbb{F}_{q}^{k}$ to $\mathbb{F}_{q}^{k}$. To specify a linear operator, we only need to define it on vectors of any basis in $\mathbb{F}_{q}^{k}$. Let $x_{1},\dots,x_{t}$ be a basis of $\Sigma_{s}(i)$. By the assumption of the lemma, vector $v_{i}$ is linearly independent with $x_{1},\dots,x_{t}$. Therefore, we can let $x_{t+1}=v_{i}$ and then extend $x_{1},\dots,x_{t},x_{t+1}$ to a basis in the space $\mathbb{F}_{q}^{k}$ with some $x_{t+2},\ldots,x_{k}$.
To define $s$, we should specify one by one linearly independent vectors $y_{1}=s.x_{1},y_{2}=s.x_{2},\ldots,y_{k}=s.x_{k}$. We have $q^{k}-1$ possibilities to choose $y_{1}$ (any non zero vector), $q^{k}-q$ possibilities to choose $y_{2}$ (any vector linearly independent with the fixed $y_{1}$), $q^{k}-q^{2}$ possibilities to choose $y_{3}$ (any vector linearly independent with $y_{1}$ and $y_{2}$), and so on. In particular, if we have fixed the values $y_{i}=s.x_{i}$ for $i=1,\ldots,t$, then it remains $q^{k}-q^{t}$ available options to choose $y_{t+1}$ (which is the same as $v_{i+1}$ in our notation).
Remark.
This implies a sort of transitivity of the group action which is stronger than the simple transitivity, but weaker than the $k$-transitivity: for all $t\leq k$, if $(x_{1},\dots,x_{t})$ and $(y_{1},\dots,y_{t})$ are two families of linearly independent elements of $\mathbb{F}_{q}^{k}$ then, there exists an element $s$ of $GL_{k}(\mathbb{F}_{q})$ such that for all $i\leq t$, $s.x_{i}=y_{i}$.
As a direct consequence of this lemma, we can formulate the following corollary:
Corollary 1.
If a matrix $s$ appears only once in $\omega$ at step $i$, $s.v_{i}=v_{i+1}$ is chosen uniformly at random among every vertex of the graph.
Proof.
Indeed, when this happens, we can rewrite $\omega$ as $AsB$ with $A$ and $B$ some invertible matrices. Then we can expose the coefficients of $A$ and $B$. The key point is that those matrices are independent from $s$, which means that we still have no information about $s$. This way, since $v_{i}=B.v_{1}$, we have $\mathbf{P}(\omega.v_{1}=v_{1})=\mathbf{P}(s.v_{i}=A^{-1}v_{1})=\frac{1}{n}$. ∎
The probability we have just found is conditioned by the structure of $\omega$, hence it needs to be multiplied by the probability of the condition. This is only dependent on the properties of $\omega$ which can be seen as an element of a free group generated by the set $\{s_{1},\dots s_{d}\}$. Let us remind that $\omega$ can be seen as a words whose letters are taken from $\{s_{1},s^{-1}_{1},\dots s_{d},s^{-1}_{d}\}$. In order to estimate the total probability of having a closed walk, we subdivide the space of such words in a few events whose probabilities will be determined. Those events are chosen so the conditional probability is more convenient to estimate. This will be done later. Let us define our events:
  • $X_{1}$ : “at least one letter appears exactly once”
  • $X_{2}$ : $\overline{X_{1}}$wedge “at least one letter appears exactly twice with same sign”
  • $X_{3}$ : “no letter appears once or twice, at least one letter appears exactly three times”
  • $X_{4}$ : “no letter appears once, twice, nor three times”
  • $X_{2}^{\prime}$ : $\overline{X_{1}}\wedge\overline{X_{3}}\wedge\overline{X_{4}}\wedge$ “all letters that appear exactly twice have different sign”
These events form a partition of the set of all possible words, whose size is $(2d)^{2m}$. In order to finish the proof of Theorem 3, let us leave aside all the events but $X_{1}$ and consider its complementary. We bound their probability using an argument similar to that of [3]. We observe that a word that belongs to $\overline{X_{1}}$ has at most $m$ different letters in it, hence we have at most $\binom{d}{m}$ ways of choosing those letters in the alphabet. When choosing each letter at random, the probability that all of them are in the right set is $(\frac{m}{d})^{2m}$. Hence
$\mathbf{P}(\overline{X_{1}})\leq\binom{d}{m}\bigg{(}\frac{m}{d}\bigg{)}^{2m}\leq\bigg{(}e\frac{d}{m}\bigg{)}^{m}\bigg{(}\frac{m}{d}\bigg{)}^{2m}=\bigg{(}e\frac{m}{d}\bigg{)}^{m}.$
The probability we are looking for is then
$\mathbf{P}(\omega.v_{1}=v_{1})=\mathbf{P}(\omega.v_{1}=v_{1}|X_{1})\mathbf{P}(X_{1})+\mathbf{P}(\omega.v_{1}=v_{1}|\overline{X_{1}})\mathbf{P}(\overline{X_{1}})$
$\leq\mathbf{P}(\omega.v_{1}=v_{1}|X_{1})+\mathbf{P}(\overline{X_{1}})\leq\frac{1}{n}+\bigg{(}e\frac{m}{d}\bigg{)}^{m}.$
We set $m=\ln n$ to minimise the bound. Substituting the above quantity back into equation 3 completes the proof of Theorem 3.
In order to find a tighter bound, we use a more careful analysis. We consider again all of our events. We are going to represent the probability $P(\omega.v_{1}=v_{1})$ as the sum
$P(\omega.v_{1}=v_{1}\mid\omega\in X_{1})\cdot P(X_{1})+\ldots+P(\omega.v_{1}=v_{1}\mid\omega\in X_{4})\cdot P(X_{4})$
${}+P(\omega.v_{1}=v_{1}\mid\omega\in X_{2}^{\prime})\cdot P(X_{2}^{\prime}).$
For $i=1,2,3$ and 4 we estimate separately $P(X_{i})$ and $P(\omega.v_{1}=v_{1}\mid\omega\in X_{i})$, and we estimate the value of the product $P(\omega.v_{1}=v_{1}\mid\omega\in X_{2}^{\prime})\cdot P(X_{2}^{\prime})$ as a whole. The sum of these bounds result in the proof of Theorem 1.
The events $X_{2}^{\prime}$ and $X_{4}$ together involve one particular event that implies a closed walk with probability $1$. This event is the “collapse” of the whole word to the identity matrix. It happens when iterating the reduction operation ($Ass^{-1}B\mapsto AB$ for all invertible matrices $A$, $s$ and $B$) ends up with the identity matrix. The probability of this event denoted $C$ is analysed in [6] (lemma 2):
$\mathbf{P}(C)=\binom{2m+1}{m}\frac{(2d)^{m}}{2m+1}\bigg{(}\frac{1}{2d}\bigg{)}^{2m}\leq\bigg{(}\frac{2}{d}\bigg{)}^{m}.$
This is proven by counting the number of well parenthesized words of size $2m$ (Catalan number) with $d$ different type of parenthesis. We wish to bound the probability of having a closed walk when $\omega$’s structure is such that this event cannot happen.
We express the size of these sets using a more general recursive formula. Let $X_{p}(c,\ell,d)$ be the size of the set of all words of length $\ell$, on the alphabet that consists of $d$ letters and its negations, such that at least $c$ letters appear (with the positive or negative sign) in this word $p$ times, and the other letters that appear in it have more occurrences. Then
$X_{p}(c,\ell,d)=\sum_{i=c}^{\lfloor\frac{\ell}{p}\rfloor}\binom{d}{i}\prod_{j=0}^{i-1}2^{p}\binom{\ell-jp}{p}X_{p+1}(0,\ell-ip,d-i)$ (4)
Indeed, $2^{p}\binom{\ell}{p}$ is the number of ways to place $p$ times the same letters in a word of size $\ell$ (each letter can have positive or negative sign). Thus, $\prod_{j=0}^{i-1}2^{p}\binom{\ell-jp}{p}$ is the number of ways of repeating $i$ times this operation while removing at every step $p$ free places. It simplifies as follows
$\prod_{j=0}^{i-1}2^{p}\binom{\ell-jp}{p}=\bigg{(}\frac{2^{p}}{p!}\bigg{)}^{i}\prod_{j=0}^{i-1}\frac{(\ell-jp)!}{(\ell-jp-p)!}=\bigg{(}\frac{2^{p}}{p!}\bigg{)}^{i}\frac{\ell!}{(\ell-ip)!}$
Then, once the $i$ different letters are placed $p$ times, we know that the $d-i$ other different letters will appear either 0 or more than $p+1$ times, and the remaining spaces will be $\ell-ip$. This explains the recursive call in 4. Note that $X_{p}(0,0,d)=1$ because we only have one way of placing no letters in a word of size 0. Moreover, if $p>\ell$ then $X_{p}(c,\ell,d)=0$, since the $p$ letters cannot fit in the word. Using these observations, we have a complete recursive definition of $X_{p}(c,\ell,d)$,
$X_{p}(c,\ell,d)=\left\{\begin{array}[]{ll}1\mbox{ if }c=0\mbox{ and }\ell=0;\\ 0\mbox{ if }p>\ell;\\ \sum_{i=c}^{\lfloor\frac{\ell}{p}\rfloor}\binom{d}{i}\bigg{(}\frac{2^{p}}{p!}\bigg{)}^{i}\frac{\ell!}{(\ell-ip)!}X_{p+1}(0,\ell-ip,d-i)\mbox{ otherwise.}\end{array}\right.$
Then we get
$|X_{1}|=X_{1}(1,2m,d),$
$|X_{3}|=X_{3}(1,2m,d)$
and
$|X_{4}|=X_{4}(0,2m,d).$
One can notice that $|X_{3}\sqcup X_{4}|=|X_{3}|+|X_{4}|=X_{3}(0,2m,d)$.
Since there are $4^{i}-2^{i}$ possibilities for choosing the sign of $i$ pairs so that the letters of at least one of them have same sign, the number of ways of placing these pairs in the word is $\binom{d}{i}(4^{i}-2^{i})\prod_{j=0}^{i-1}\binom{2m-2j}{2}=\binom{d}{i}(2^{i}-1)\frac{(2m)!}{(2m-2i)!}$. Hence
$|X_{2}|=\sum_{i=1}^{m}\binom{d}{i}(2^{i}-1)\frac{(2m)!}{(2m-2i)!}X_{3}(0,2m-2i,d-i).$
Similarly, there are $2^{i}$ ways of choosing the sign of $i$ pairs of letter so that all pairs are of different sign. Thus we get
$|X_{2}^{\prime}|=\sum_{i=1}^{m}\binom{d}{i}\frac{(2m)!}{(2m-2i)!}X_{3}(0,2m-2i,d-i).$
This can be summarized with the relation $X_{2}(1,2m,d)=|X_{2}|+|X_{2}^{\prime}|$.
We have already explained that $\mathbf{P}(\omega.v_{1}=v_{1}|X_{1})=\frac{1}{n}$. It remains to bound this probability conditioned to the other events. In what is next, we will set $q=2$. Taking a bigger field might weaken the bounds and complicate the analysis. We start with $X_{2}$.
Lemma A.1.2.
$\mathbf{P}(\omega.v_{1}=v_{1}|X_{2})\leq\frac{2}{n}.$
Proof.
Let $s$ be the matrix that appears twice with same sign. The word is then of the form $AsBsC.v_{1}=v_{1}$ with $A,B$ and $C$ some invertible matrices of known coefficients. We can rewrite this equation as $sBs.x=y$ with $x$ and $y$ two determined vectors ($x=C.v_{1}$ and $y=A^{-1}.v_{1}$). It is useful to name the different vectors of the product:
$\underbrace{s\ \overbrace{B\ \underbrace{s.x}_{y^{\prime}}}^{x^{\prime}}}_{y^{\prime\prime}}=y.$
Since we are in the field of size two there is no non-trivial pairs of parallel vectors. Hence the step that leads to $y^{\prime\prime}$ is free only if $x^{\prime}\neq x$. In a larger field ($q>2$), for $y^{\prime\prime}$ to be free, it is necessary that $x\neq\alpha x^{\prime}$ for all non zero $\alpha\in\mathbb{F}_{q}$. By taking $q=2$, a lot of case-by-case analysis is avoided.
Because $y^{\prime}$ is necessarily free and since $x$ and $x^{\prime}=B.y^{\prime}$ are independent, $\mathbf{P}(x^{\prime}=x)=\frac{1}{n}$. Then, if $x^{\prime}=x$, we have $y^{\prime\prime}=y^{\prime}$. This is the probability that $y^{\prime\prime}$ is forced. If this is not the case, namely if $x^{\prime}\neq x$ (which happens with probability $\frac{n-1}{n}$), then the probability for $y^{\prime\prime}$ to be equal to $y$ is at most $\frac{1}{n-1}$ ($y^{\prime\prime}$ cannot be equal to $y^{\prime}$ since both steps are free). Therefore,
$\mathbf{P}(\omega.v_{1}=v_{1}|X_{2})\leq\frac{1}{n}+\frac{n-1}{n}\frac{1}{n-1}=\frac{2}{n}.$
Now we bound the probability $\mathbf{P}(\omega v_{1}=v_{1}|X_{3})$. We proceed the same way as above, by distinguishing the cases where the final step is free or not. We prove the following claim:
Lemma A.1.3.
$\mathbf{P}(\omega.v_{1}=v_{1}|``at\ least\ a\ letter\ appears\ exactly\ three\ times")\leq\frac{5}{n}.$
In particular, we have
$\mathbf{P}(\omega.v_{1}=v_{1}|X_{3})\leq\frac{5}{n}.$
Proof.
Under the $X_{3}$ condition, the word can take four forms that will be analysed separately:
  • $sBsCs.x=y$
  • $s^{-1}BsCs.x=y$
  • $sBs^{-1}Cs.x=y$
  • $sBsCs^{-1}.x=y$
Any other form can be turned into one of the above by switching $s$ with $s^{-1}$, which does not change the argument. The different possibilities can be summarized by writing $s_{1}Bs_{2}Cs_{3}.x=y$; at most one of $s_{1},s_{2},s_{3}$ is $s^{-1}$ and the others are $s$. We will use the notation below to treat all four cases:
$\underbrace{s_{1}\ \overbrace{B\ \underbrace{s_{2}\ \overbrace{C\ \underbrace{s_{3}.x}_{y^{\prime}}}^{x^{\prime}}}_{y^{\prime\prime}}}^{x^{\prime\prime}}}_{y^{\prime\prime\prime}}=y.$
We start with the case in which there is no $s^{-1}$ in $\omega$. Here, the step that leads to $y^{\prime\prime\prime}$ is free only if $x^{\prime\prime}\neq x$, $x^{\prime\prime}\neq x^{\prime}$ and $x^{\prime\prime}\neq x^{\prime}+x^{\prime\prime}$ (that is, $x^{\prime\prime}$ is not a linear combination of $x^{\prime}$ and $x$). Since $x^{\prime}$ and $x$ are independent, $\mathbf{P}(x^{\prime}=x)=\frac{1}{n}$. Hence, with probability $\frac{n-1}{n}$ we get that $y^{\prime\prime}$ is free, which means that $x^{\prime\prime}=B.y^{\prime\prime}$ is uniformly distributed among the $n-1$ vectors different from $B.y^{\prime}$. There are three values for $x^{\prime\prime}$ that make the final step forced and they are equally likely, thus $\mathbf{P}(y^{\prime\prime\prime}\text{ is forced }|x^{\prime}\neq x)\leq\frac{3}{n-1}$. The opposite case happens with probability $\frac{n-4}{n-1}$. Then $\mathbf{P}(y^{\prime\prime\prime}=y)\leq\frac{1}{n-3}$. To illustrate the reasoning, we can represent those probabilities by a tree:
$x^{\prime}=x$$x^{\prime}\neq x$$x^{\prime\prime}=x$$x^{\prime\prime}\neq x$$x^{\prime\prime}=x^{\prime}$$x^{\prime\prime}\neq x^{\prime}$$x^{\prime\prime}=x+x^{\prime}$$x^{\prime\prime}\neq x+x^{\prime}$$\frac{1}{n}$$\frac{n-1}{n}$$\frac{1}{n-1}$$\frac{n-2}{n-1}$$\frac{1}{n-2}$$\frac{n-3}{n-2}$$\frac{1}{n-3}$$\frac{n-4}{n-3}$
The rightmost leaf corresponds to $y^{\prime\prime\prime}$ being free which gives a probability $\frac{1}{n-3}$ of having a closed walk. Therefore,
$\mathbf{P}(sBsCs.x=y)\leq\frac{1}{n}+\frac{n-1}{n}\bigg{(}\frac{3}{n-1}+\frac{n-4}{n-1}\frac{1}{n-3}\bigg{)}\leq\frac{5}{n}.$
Now, consider $s_{3}=s^{-1}$. Then $y^{\prime\prime}$ is forced if $x^{\prime}=y^{\prime}$, but those two vectors are correlated, so we cannot bound the probability of this event. We will consider both cases and take the probability of the most likely event as a bound. If $y^{\prime}=x^{\prime}$, then $y^{\prime\prime}$ is forced, which implies that $y^{\prime\prime}=x$. In this case, if $x^{\prime\prime}=y^{\prime}$ we have $y^{\prime\prime\prime}=x$. However, $x^{\prime\prime}=B.x$, which is independent from $y^{\prime}$ (which is from a free step). Hence, the probability for them to be equal is $\frac{1}{n}.$ In the opposite case, $y^{\prime\prime\prime}$ is free, which gives a total probability of this branch of $\frac{2}{n}$. We now suppose that $y^{\prime\prime}$ is free. Then, with probability $\frac{3}{n-1}$, $y^{\prime\prime\prime}$ is forced. In the other case, $y^{\prime\prime\prime}$ is equal to $y$ with probability at most $\frac{1}{n-3}$. Here is the probability tree:
$x^{\prime}=y^{\prime}$$x^{\prime\prime}\neq y^{\prime}$$x^{\prime\prime}=y^{\prime}$$x^{\prime}\neq y^{\prime}$$x^{\prime\prime}=x^{\prime}$$x^{\prime\prime}\neq x^{\prime}$$x^{\prime\prime}=y^{\prime}$$x^{\prime\prime}\neq y^{\prime}$$x^{\prime\prime}=y^{\prime}+x^{\prime}$$x^{\prime\prime}\neq y^{\prime}+x^{\prime}$$\frac{n-1}{n}$$\frac{1}{n}$$\frac{1}{n-1}$$\frac{n-2}{n-1}$$\frac{1}{n-2}$$\frac{n-3}{n-2}$$\frac{1}{n-3}$$\frac{n-4}{n-3}$
The branches whose probabilities are close to 1 corresponds to the cases when $y^{\prime\prime\prime}$ is free. Thus we have
$\mathbf{P}(sBsCs^{-1}.x=y)\leq\max\bigg{(}\frac{2}{n},\frac{3}{n-1}+\frac{n-4}{n-1}\frac{1}{n-3}\bigg{)}\leq\frac{4}{n-1}.$
When $s_{2}=s^{-1}$ the choice of $y^{\prime\prime}$ is forced if $x^{\prime}=y^{\prime}$, which implies $x^{\prime\prime}=B.x$. Then $s$ is defined only on $x$. If $B.x=x$ then, since $y^{\prime}$ is free, $\mathbf{P}(y^{\prime\prime\prime}=y)=\mathbf{P}(y^{\prime}=y)=\frac{1}{n}$. Otherwise, $y^{\prime\prime\prime}$ is free, and therefore $\mathbf{P}(y^{\prime\prime\prime}=y)=\frac{1}{n-1}$. On the other hand, if $y^{\prime\prime}$ is free, then $s$ is defined on $x$ and $y^{\prime\prime}$. Since $y^{\prime\prime}$ is free, $x^{\prime\prime}$ and $x$ are independent, thus $\mathbf{P}(x^{\prime\prime}=x)=\frac{1}{n-1}$. Moreover, if $x^{\prime\prime}=y^{\prime\prime}$, $\mathbf{P}(y^{\prime\prime\prime}=y)=\mathbf{P}(x^{\prime}=y)=\frac{1}{n}$. If $x^{\prime\prime}=y^{\prime\prime}+x$, we have $y^{\prime\prime\prime}=x^{\prime}+y^{\prime}=(C+Id_{k})y^{\prime}$ which is independent from $y$. Hence, $\mathbf{P}(y^{\prime\prime\prime}=y)=\frac{1}{n}$. Otherwise, since three vectors are excluded, $y^{\prime\prime\prime}$ is free with probability at most $\frac{n-3}{n}$. If so, $\mathbf{P}(y^{\prime\prime}=y)=\frac{1}{n-3}$. As before, the case study can be illustrated with a tree:
$x^{\prime}=y^{\prime}$$x^{\prime\prime}\neq x$$x^{\prime\prime}=x$$x^{\prime}\neq y^{\prime}$$x^{\prime\prime}=x$$x^{\prime\prime}\neq x$$x^{\prime\prime}=y^{\prime\prime}$$x^{\prime\prime}\neq y^{\prime\prime}$$x^{\prime\prime}=y^{\prime\prime}+x$$x^{\prime\prime}\neq y^{\prime\prime}+x$$\frac{1}{n-1}$$\frac{n-2}{n-1}$$\frac{1}{n-2}$$\frac{n-3}{n-2}$
Thus, we get
$\mathbf{P}(sBs^{-1}Cs.x=y)\leq\max\bigg{(}\frac{2}{n-1},\frac{3}{n-1}+\frac{n-2}{n-1}\bigg{(}\frac{n-3}{n-2}\frac{1}{n-3}\bigg{)}\bigg{)}=\frac{4}{n-1}.$
Lastly, we consider the case $s_{1}=s^{-1}$. Here, $y^{\prime\prime}$ is forced when $x^{\prime}=x$. Those are not correlated, so this event happens with the probability $\frac{1}{n}$. In the other case, $s^{-1}$ is defined on $y^{\prime\prime}$ and $y^{\prime}$, which are random. If $x^{\prime\prime}=y^{\prime\prime}$, we have $y^{\prime\prime\prime}=x^{\prime}$ which is equal to $y$ with probability less than $\frac{1}{n-1}$. Since $y^{\prime\prime}$ is free, $y^{\prime}$ and $y^{\prime\prime}$ are independent, hence $\mathbf{P}(x^{\prime\prime}=y^{\prime})=\mathbf{P}(x^{\prime\prime}=y^{\prime}+y^{\prime\prime})=\frac{1}{n-1}$. If $y^{\prime\prime\prime}$ is free, it can take any value with probability $\frac{1}{n-3}$. The last probability tree is then
$x^{\prime}=x$$x^{\prime}\neq x$$x^{\prime\prime}=y^{\prime\prime}$$x^{\prime\prime}\neq y^{\prime\prime}$$x^{\prime\prime}=y^{\prime}$$x^{\prime\prime}\neq y^{\prime}$$x^{\prime\prime}=y^{\prime}+y^{\prime\prime}$$x^{\prime\prime}\neq y^{\prime}+y^{\prime\prime}$$\frac{1}{n}$$\frac{n-1}{n}$$\frac{1}{n-2}$$\frac{n-3}{n-2}$$\frac{1}{n-3}$$\frac{n-4}{n-3}$
Hence we have
$\mathbf{P}(s^{-1}BsCs.x=1)\leq\frac{1}{n}+\frac{n-1}{n}\max\bigg{(}\frac{1}{n-2},\frac{1}{n-2}+\frac{n-4}{n-2}\frac{1}{n-3}\bigg{)}\leq\frac{5}{n}.$
By taking the maximum of all these bounds, we conclude the proof. ∎
It remains to bound $\mathbf{P}(\omega.v_{1}=v_{1}|X_{2}^{\prime})\mathbf{P}(X_{2}^{\prime})$. To simplify the notations we set $x_{3}(i)=X_{3}(1,2m-2i,d-i)$ and $x_{4}(i)=X_{4}(0,2m-2i,d-i)$. We need to prove the following statement.
Lemma A.1.4.
$\mathbf{P}(\omega.v_{1}=v_{1}|X_{2}^{\prime})\mathbf{P}(X_{2}^{\prime})\leq\\ \bigg{(}\frac{1}{2d}\bigg{)}^{2m}\sum_{i=1}^{m}\binom{d}{i}\frac{(2m)!}{(2m-2i)!}\bigg{[}\frac{x_{3}(i)+x_{4}(i)}{n}+\frac{2^{i}}{(i+1)!}\bigg{(}\frac{5}{n}x_{3}(i)+x_{4}(i)\bigg{)}\bigg{]}.$ (5)
Proof.
Before we proceed with the proof of this lemma we stress again that in this statement we do not bound separately $\mathbf{P}(X_{2}^{\prime})$ and $\mathbf{P}(\omega.v_{1}=v_{1}|X_{2}^{\prime})$, we estimate directly the product of these two probabilities, which equals to the probability of the event
$\mathbf{P}(\omega.v_{1}=v_{1}\text{ and }\omega\in X_{2}^{\prime}).$
The probability is taken, as usual, over the random choice of a word $\omega$ of $2m$ letters and the random choice of invertible matrices assigned to the letters of this alphabet.
We start the proof with two claims.
Claim 1: Assume that the word $\omega$ contains letters $t$ and $s$ exactly twice, and each of these letters appears once with the positive and once with the negative sign, and these letters interleave:
$\omega=\ldots t\ldots s\ldots t^{-1}\ldots s^{-1}\ldots$ (6)
Then the probability to get a closed walk corresponding to the path $\omega$ (probability taken over the choice of matrices for each letter in the alphabet) is equal to $1/n$. The claim remains true if we swap the positions of the pair of letters $s$ and $s^{-1}$ and/or of the pair of letters $t$ and $t^{-1}$.
Proof of the claim.
Words from $X_{2}^{\prime}$ are all of the form $AsBs^{-1}C$ with $A$, $B$ and $C$ some invertible matrices. Hence, we wish to estimate the probability of the event $sBs^{-1}.x=y$, with $x=C.v_{1}$, and $y=A^{-1}.v_{1}$. We use the notation
$\underbrace{s\ \overbrace{B\ \underbrace{s^{-1}.x}_{y^{\prime}}}^{x^{\prime}}}_{y^{\prime\prime}}=y.$
Here, the matrix $t$ is a factor of $B$ (hence $B=\dots t\dots$). We first suppose that $x=y$. Then, if $x^{\prime}=y^{\prime}$ we have $y^{\prime\prime}=x=y$. Since $t$ appears in $B$, $x^{\prime}$ is independent of $y^{\prime}$, and thus $\mathbf{P}(y^{\prime}=x^{\prime})=\frac{1}{n}$ (because this is the first time $t$ is used in the path). In the opposite case, we have $y^{\prime\prime}\neq y$, and the path cannot be closed.
Now we suppose $x\not=y$. Then if $y^{\prime}=x^{\prime}$ we have $y^{\prime\prime}=x\neq y$. If $y^{\prime}\neq x^{\prime}$ (which happens with probability $\frac{n-1}{n}$, $y^{\prime\prime}$ is free, and its value is uniformly distributed among the $n-1$ remaining vectors. Therefore, when we have this configuration of random matrices in $\omega$, the probability of having a closed walk is $\frac{1}{n}$. ∎
It can be noticed that here, the fact that $t$ appears with different sign is not used.
Claim 2: Let us take the set of $2i$ literals
$\{s_{1},s_{1}^{-1},\ldots,s_{i},s_{i}^{-1}\}$
and consider the set of all words of length $(2i)$ composed of these literals (each one should be used exactly once). We claim that the fraction of words that represent a well formed structure of $i$ pairs of parentheses, where each pairs is associated with some pair of literals $(s_{j},s^{-1}_{j})$ or $(s^{-1}_{j},s_{j})$, is equal to $\frac{2^{i}}{(i+1)!}$.
Proof of claim.
In general, we have $(2i)!$ different ways to distribute $(2i)$ literals among $(2i)$ positions. Let us count the fraction of permutations where the literals form a structure of $i$ pairs of parentheses. The number of well parenthesized words (with one type of parentheses) of size $2i$ is the Catalan number $\binom{2i+1}{i}\frac{1}{2i+1}$. We have $i!$ ways to assign each pair of parentheses with one of $i$ types of literals, and $2^{i}$ to chose the signs in each pairs ($\ldots s_{j}\ldots s_{j}^{-1}\ldots$ or $\ldots s_{j}^{-1}\ldots s_{j}\ldots$ for each of $i$ pairs). Hence, the proportion of the well parenthesized words is
$\frac{\binom{2i+1}{i}\frac{1}{2i+1}i!2^{i}}{(2i)!}=\frac{(2i+1)!i!2^{i}}{(2i+1)!i!(i+1)!}=\frac{2^{i}}{(i+1)!}.$
It is easy to see that the absence of pattern 6 is equivalent to having such well formed structure of parenthesis.
Let us proceed with the proof of the lemma. By definition, in each word $\omega\in X_{2}^{\prime}$ all letters that appear exactly twice must have different signs. In what follows we denote $i$ that number of letters that appear in $\omega$ exactly twice. For a fixed $i$, to specify a word $\omega$ where $i$ letters appear twice (with opposite signs) and the other letters appear at least three times, we should
  • choose $i$ letters among $d$ (those who appear exactly twice), which gives $\binom{d}{i}$ combinations;
  • choose $2i$ positions in the word $\omega$ of length $2m$ where we place the letters that appear twice, which gives $\binom{2m}{2i}$ combination;
  • fix a permutations of the $(2i)$ literals on the chosen $(2i)$ positions, which gives $(2i)!$ combinations;
  • fill the remaining $(2m-2i)$ positions of $\omega$ with other letters, using each letter at least three times; we subdivide these combinations into two subcases:
    • there is at least one letter that is used exactly three times; we have $x_{3}(i)=X_{3}(1,2m-2i,d-i)$ possibilities to do it;
    • there is no letter that is used exactly three times, i.e., each letter (besides the $i$ letters that were used twice) must be used at least four times; we have $x_{4}(i)=X_{4}(0,2m-2i,d-i)$ possibilities to fill in this way the remaining $(2m-2i)$ positions.
The $i$ pairs of letters in $\omega$ contain the pattern (6) may contain or not contain the pattern (6). By Claim 2, the latter is the case for the fraction $\frac{2^{i}}{(i+1)!}$ of all $\omega$ (with $i$ pairs) and, respectively, the former is the case for the fraction $1-\frac{2^{i}}{(i+1)!}$ of these words.
If the $i$ pairs of letters in $\omega$ contain the pattern (6), then by Claim 1 the probability that $\omega$ provides a closed path is at most $\frac{1}{n}$ (probability taken over the choice of matrices for each letter in the alphabet). Since we have in total $(2d)^{2m}$ words $\omega$, this case contributes to the resulting probability $\mathbf{P}(\omega.v_{1}=v_{1}\text{ and }\omega\in X_{2}^{\prime})$ at most
$\bigg{(}\frac{1}{2d}\bigg{)}^{2m}\binom{d}{i}\binom{2m}{2i}(2i)!\bigg{(}1-\frac{2^{i}}{(i+1)!}\bigg{)}(x_{3}(i)+x_{4}(i))\cdot\frac{1}{n}$
(in what follows we bound $1-\frac{2^{i}}{(i+1)!}$ by $1$).
If the $i$ pairs of letters in $\omega$ do not contain the pattern (6) but one of other letters appear in $\omega$ exactly thee times, then the probability to have a closed path is at most $\frac{5}{n}$, as shown in Lemma A.1.3. This case contributes to the resulting probability at most
$\bigg{(}\frac{1}{2d}\bigg{)}^{2m}\binom{d}{i}\binom{2m}{2i}(2i)!\cdot\frac{2^{i}}{(i+1)!}\cdot x_{3}(i)\cdot\frac{5}{n}$
At last, if $\omega$ does not contain the pattern (6) and all other letters appearing in $\omega$ are used more than three times, then we trivially bound the probability to have a closed path by $1$. This contributes to the resulting probability
$\bigg{(}\frac{1}{2d}\bigg{)}^{2m}\binom{d}{i}\binom{2m}{2i}(2i)!\cdot\frac{2^{i}}{(i+1)!}\cdot x_{4}(i).$
Summing these quantities for all possible values of $i$ and observing that $\binom{2m}{2i}(2i)!=\frac{(2m)!}{(2m-2i)!}$, we obtain the statement of the lemma.
We proceed with similar bounds for the other sets of words:
$\mathbf{P}(\omega.v_{1}=v_{1}|X_{1})\mathbf{P}(X_{1})\leq\frac{1}{n}\bigg{(}\frac{1}{2d}\bigg{)}^{2m}|X_{1}|,$
$\mathbf{P}(\omega.v_{1}=v_{1}|X_{2})\mathbf{P}(X_{2})\leq\frac{2}{n}\bigg{(}\frac{1}{2d}\bigg{)}^{2m}|X_{2}|,$
$\mathbf{P}(\omega.v_{1}=v_{1}|X_{3})\mathbf{P}(X_{3})\leq\frac{5}{n}\bigg{(}\frac{1}{2d}\bigg{)}^{2m}|X_{2}|,$
and
$\mathbf{P}(\omega.v_{1}=v_{1}|X_{4})\mathbf{P}(X_{4})\leq\bigg{(}\frac{1}{2d}\bigg{)}^{2m}|X_{4}|.$
The sum of these expressions is larger than $P_{11}$ defined at the beginning of this section. By replacing it in equation 3, we complete the proof. It is easy to see that the rough bounding used in the proof of Theorem 3 gives a larger bound than that of Theorem 2.
A.2 Proof of Theorem 2
[…]
$M=\left(\begin{array}[]{@{}c|c@{}}0&A\\ \hline\cr\,{}^{t}A&0\end{array}\right)$
where ${}^{t}A$ is the transposition of $A$. In a bipartite graph, it is known that the spectrum $|\mu_{1}|\geq\dots\geq|\mu_{2n}|$ is symmetric with respect to zero. Hence for $1\leq i\leq n$, we have $|\mu_{2i+1}|=|\mu_{2(i+1)}|$. This way we get
$\sum_{i=0}^{n-1}2\mu_{2i+1}^{2m}=Trace(M^{2m}).$
In order to study the spectral gap, the relevant quantity to estimate is then $|\mu_{3}|=|\mu_{4}|$. Since $|\mu_{1}|=|\mu_{2}|=1$, we thus obtain
$22\mu_{3}^{2m}\leq\sum_{i=1}^{n-1}2\mu_{2i+1}^{2m}=Trace(M^{2m})-$ 2 .
As we have seen in section 3.1, the expected value of $Trace(M^{2m})$ is the sum of the probability of getting of closed path of size $2m$, starting on each vertex. We note this probability $P_{ii}$ for the vertex $i$. It is the same for every vertex, hence we get, by using Jensen’s inequality
$\mathbf{E}(|\mu_{3}|)\leq\bigg{(}\frac{1}{2}(\mathbf{E}(Trace(M^{2m}))-2)\bigg{)}^{\frac{1}{2m}}=(nP_{11}-1)^{\frac{1}{2m}}.$ (7)
One can notice that here, $n$ is the size of the partition, not the size of the graph. Indeed, with an even number of steps, the path must end in the same partition as it started, which eliminates half of the vertices.
We first explain why the construction for even degree regular bipartite graphs gives the same bound as Theorem 1. Here, a random walk can be represented as a sequence of matrices of $D\cup D^{-1}$. This is because every vertex $x$ is connected to $s.x$ and $s^{-1}.x$. Each element of the sequence is chosen independently of the others. It is then easy to see that the structure of the walk is exactly the same as in the non bipartite case: a uniformly random sequence of $2m$ matrices from $D\cup D^{-1}$. The same proof can then be applied to this sequence, the elements of the sequence will then behave the same way as in the preceding section.
However, some work needs to be done for graphs of odd degree. In order to apply here a similar reasoning as in the previous section, we need to understand what a random walk in $G$ looks like in terms of the matrices of $D$.
[…]
Figure 6: The steps of the random walk work by pairs of matrices of $D$; the first one brings us on a vertex of the right hand side, the other is for the way back.
In a bipartite regular graph obtained by our construction, a random walk of size $2m$ is a sequence $\omega$ of elements of $D$ chosen independently at random. As usual, every edge in the graph corresponds to some invertible matrix. In the case of a bipartite graph we assume that the multiplication by the matrices transforms the vertices in the left part into the vertices of the right part. Thus, in a random walk on such a graph, the matrices that appear in an odd position in $\omega$ are taken with the positive sign, while the matrices that appear in an even step are taken with the negative sign, see Fig. 6.
We wish to proceed as in the preceding section, by partitioning the set of possible sequences (words) so that we can analyse the probability of having a closed walk conditioned to the sets this partition. In order to reuse the results above, we choose a similar partitioning. There is a minor difference: now the signs of matrices appearing in $\omega$ are fixed (positive for odd position and negative for the others). Hence, the number of possible words is $d^{2m}$. We define the sets before determining their size:
  • $Y_{1}$ : “at least one letter appears exactly once”
  • $Y_{2}$ : $\overline{Y_{1}}$wedge ”at least one letter appears exactly twice”
  • $Y_{3}$ : “no letter appear once or twice, at least one letter appears exactly three times”
  • $Y_{4}$ : “no letter appear once, twice, nor three times”
Up to this point, it is not hard to understand why the bound of Theorem 3 holds. Indeed, the sign of the matrices do not play any role in the proof, so the probability of $\overline{Y_{1}}$ can be bounded by the same quantity as in the previous section. In addition, the probability of having a closed walk conditioned to $Y_{1}$ is also $\frac{1}{n}$ ($n$ is the size of a partition). Therefore, Theorem 3 applies to bipartite regular graphs.
We can define the analogous recursive relation used in the preceding part. Since this formula represents a quantity that does not depend on the sign of the letters (they are determined by the parity of the positions), we can just ignore them:
$Y_{p}(c,\ell,d)=\left\{\begin{array}[]{ll}1\mbox{ if }c=0\mbox{ and }\ell=0\\ 0\mbox{ if }p>\ell\\ \sum_{i=c}^{\lfloor\frac{\ell}{p}\rfloor}\binom{d}{i}(p!)^{-i}\frac{\ell!}{(\ell-ip)!}Y_{p+1}(0,\ell-ip,d-i)\mbox{ otherwise.}\end{array}\right.$
As before, $Y_{p}(c,\ell,d)$ is the number of words of size $\ell$ on alphabet of size $d$ that have at least $c$ different letters that appear $p$ times and whose other present letters have more occurrences. The only difference with $X_{p}(c,\ell,d)$ is that we do not deal with signs. For the same reason, we have
$|Y_{1}|=Y_{1}(1,2m,d),$
$|Y_{2}|=Y_{2}(1,2m,d),$
$|Y_{3}|=Y_{3}(1,2m,d)$
and
$|Y_{4}|=Y_{4}(0,2m,d).$
We have already shown that when one letter appears exactly once, the probability of having a closed walk is $\frac{1}{n}$. Similarly, when a letter appears exactly three times, the probability of getting a closed walk is less than $\frac{5}{n}$. Indeed, in the proof of Lemma A.1.3, all possible configurations of signs for the letter that appears three times are considered (e.g. $\omega=\dots s\dots s^{-1}\dots s\dots$ or $\dots s\dots s\dots s^{-1}\dots$). No assumption is done on their respective probabilities to occur. These probabilities may or may not be different in the bipartite setting. Since this bound ($\frac{5}{n}$) is the maximum over all the probabilities of getting a closed walk with each configuration of signs, the resulting bound for the probability of getting a closed walk in the bipartite case remains the same. Hence we get
$\mathbf{P}(\omega.v_{1}=v_{1}|Y_{1})\mathbf{P}(Y_{1})\leq\bigg{(}\frac{1}{d}\bigg{)}^{2m}Y_{1}(1,2m,d)\frac{1}{n}$
and
$\mathbf{P}(\omega.v_{1}=v_{1}|Y_{3})\mathbf{P}(Y_{3})\leq\bigg{(}\frac{1}{d}\bigg{)}^{2m}Y_{3}(1,2m,d)\frac{5}{n}.$
As before, we do not bound the probability of getting a closed walk under condition $Y_{4}$. Then
$\mathbf{P}(\omega.v_{1}=v_{1}|Y_{4})\mathbf{P}(Y_{4})\leq\bigg{(}\frac{1}{d}\bigg{)}^{2m}Y_{4}(0,2m,d).$
We now estimate $\mathbf{P}(\omega.v_{1}=v_{1}|Y_{2})\mathbf{P}(Y_{2})$. We set $y_{3}(i)=Y_{3}(1,2m-2i,d-i)$ and $y_{4}(i)=Y_{4}(0,2m-2i,d-i)$.
Lemma A.2.1.
$\mathbf{P}(\omega.v_{1}=v_{1}|Y_{2})\mathbf{P}(Y_{2})\leq\\ Y_{2}(1,2m,d)\frac{2}{n}+\sum_{i=1}^{m}\binom{d}{i}\bigg{(}\frac{1}{2}\bigg{)}^{i}\frac{(2m!)}{(2m-2i)!}\frac{2^{i}}{(i+1)!}\bigg{(}y_{3}(i)\frac{5}{n}+y_{4}(i)\bigg{)}$ (8)
Proof.
We proceed in a similar way as in the proof of lemma A.1.4. Consider we have $i$ pairs of matrices that appear exactly twice in $\omega$. In the bipartite setting, the signs are forced by the parity of the position of each letter. We thus choose to ignore them. Then, the probability of having no pair $(s,t)$ such that we have the pattern $\omega=\dots s\dots t\dots s\dots t\dots$ is $\frac{2^{i}}{(i+1)!}$. The proof is the same as that of claim 2 in Lemma A.1.4, except that the numerator and the denominator of the fraction are both divided by $2^{i}$ (because we ignore the signs).
Using the proof of Lemma A.1.2, we conclude that, if a letter appears twice with the same sign, the probability of having a closed walk is less than $\frac{2}{n}$. If there is a pair $(s,t),\omega=\dots s\dots t\dots s^{-1}\dots t^{-1}\dots$, then, by using the argument from lemma A.1.4 (claim 1), the probability of getting a closed walk is $\frac{1}{n}$. If those cases do not happen, we still can bound the probability of getting a closed walk using lemma A.1.3 when at least a letter appears three times. The conditional probability of getting a closed walk is then less than $\frac{5}{n}$.
Let us combine together all these bounds.
  • The union of the event in which a letter appears exactly twice with same sign, and the event where the letters that appear twice form a bad parenthesized word (if we forget about the signs) has size smaller than $Y_{2}(1,2m,d)$. The probability of getting a closed walk in this case is not greater than $\frac{2}{n}$.
  • The size of the event in which this bound does not apply, but we can apply the bound from lemma A.1.3 (which is $\frac{5}{n}$) can be computed as follows. $(2i)!\binom{2m}{2i}$ is the number of ways of placing $i$ pairs of letters with different sign in $2m$ positions. Since we ignore the sign, this quantity has to be divided by $2^{i}$ which gives $(\frac{1}{2})^{i}\frac{(2m!)}{(2m-2i)!}$. A fraction $\frac{2^{i}}{(i+1)!}$ of them are well formed. $y_{3}(i)$ is the numbers of ways of filling the remaining gaps so that no letter appear once nor twice, and at least letter appears three times. Choosing the $i$ pairs among the $d$ possible ones and summing over all $i\leq 2m$, we get
    $\sum_{i=1}^{m}\binom{d}{i}\bigg{(}\frac{1}{2}\bigg{)}^{i}\frac{(2m!)}{(2m-2i)!}\frac{2^{i}}{(i+1)!}y_{3}(i).$
  • Similarly, the number of remaining words that correspond to walks whose probability of being closed is not estimated is
    $\sum_{i=1}^{m}\binom{d}{i}\bigg{(}\frac{1}{2}\bigg{)}^{i}\frac{(2m!)}{(2m-2i)!}\frac{2^{i}}{(i+1)!}y_{4}(i).$
Summing all theses quantities, multiplying them by their respective probabilities of getting a closed walk and dividing the whole expression by the number of possible words (that is $d^{2m}$), we can conclude.
Summing all the above probabilities and substituting this in 7 finishes the proof of Theorem 2.
A.3 Proof of Proposition 1
We now turn to bound the second largest eigenvalue for biregular graphs. Let $G^{\prime}$ be a bipartite regular graph of degree $d_{1}$ and whose number of vertices is $2n_{1}$. Let $\gamma$ be an integer that divises $n_{1}$. We construct the $d_{1}d_{2}$-biregular bipartite graph $G$ by merging every $\gamma$ vertices in the right partition. We denote $n_{2}=\frac{n_{1}}{\gamma}$ the size of this partition, and $d_{2}=\gamma d_{1}$, thus $d_{1}$ and $d_{2}$ are the respective degrees of each partition of the graph. Let
$P=\left(\begin{array}[]{@{}c|c@{}}0&M\\ \hline\cr\,{}^{t}M&0\end{array}\right)$
be its adjacency matrix. Hence $M$ has dimension $n_{1}\times n_{2}$. Let
$Q=\left(\begin{array}[]{@{}c|c@{}}0&A\\ \hline\cr\,{}^{t}A&0\end{array}\right)$
be the adjacency matrix of $G^{\prime}$, which is the bipartite regular graph before merging the vertices of the right partition. We set $J$ such that $M.^{t}M=A.J.^{t}A$, thus $J=I_{n_{2}}\otimes J_{\gamma}$, where $J_{\gamma}$ is the $\gamma\times\gamma$ matrix whose entries are only ones and $\otimes$ is the Kronecker product. An example of resulting paths is shown in Fig. 7.
${}^{t}A$$J$$A$
Figure 7: Traveling three steps starting from the left in this directed graph is equivalent to doing two steps in the bipartite graph (starting from the left as well) with merged vertices on the right. The merging operation is represented by $J$ which corresponds to the complete bipartite graphs in the middle. Here $n_{1}=6,n_{2}=2,d_{1}=1,d_{2}=3$.
We set $A=(a_{ij})_{i,j\in[\![1,\ n_{1}]\!]}$. All $A$’s columns and rows sum up to $d_{1}$ —so does ${}^{t}A$. We can show that for every $x=(x_{1},\dots,x_{n_{1}})$ orthogonal to $e_{1}=\frac{1}{\sqrt{n_{1}}}(1,\dots,1)$, $Ax$ is also orthogonal to $e$. Indeed, the coordinates of such an $x$ sum up to zero. We denote $Ax=(y_{1},\dots,y_{n_{1}})$. Then
$\sum_{i=1}^{n_{1}}y_{i}=\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{1}}a_{ij}x_{j}=\sum_{j=1}^{n_{1}}x_{j}\sum_{i=1}^{n_{1}}a_{ij}=d_{1}\sum_{j=1}^{n_{1}}x_{j}=0.$
Therefore, $Ax$ is orthogonal to $e$. The same is true for ${}^{t}Ax$.
On the other hand, it is easy to see that the spectrum of $J$ is
$\underbrace{(\gamma,...,\gamma,}_{n_{2}}\underbrace{0,...,0)}_{n_{1}-n_{2}}.$
Let $\lambda_{2}(X)$ be the second largest eigenvalue of some square matrix $X$ and let $x$ be the normalised eigenvector associated to $\lambda_{2}(M.^{t}M)$. For every positive number $q$, we have
$|\lambda_{2}(M.^{t}M)^{q}|=||(A.J.^{t}A)^{q}.x||=||A(J.^{t}A.A)^{q-1}J.^{t}A.x||\leq\gamma d_{1}||A(J.^{t}A.A)^{q-1}x^{\prime}||$
with $x^{\prime}$ a normalised vector orthogonal to $e$ (because of the preceding fact). Hence
$||(J.^{t}A.A)^{q-1}x^{\prime}||\leq\gamma^{q-1}|\lambda_{2}^{q-1}(^{t}A.A)|.$
We conclude that
$|\lambda_{2}(M.^{t}M)^{q}|\leq d_{1}^{2}\gamma^{q}|\lambda_{2}(^{t}A.A)^{q-1}|=d_{1}d_{2}(\gamma|\lambda_{2}(^{t}A.A)|)^{q-1}.$
Since this is a positive quantity, its $q$-th root is defined:
$\lambda_{2}(M.^{t}M)\leq\bigg{(}d_{1}d_{2}(\gamma|\lambda_{2}(^{t}A.A)|)^{q-1}\bigg{)}^{\frac{1}{q}}=\bigg{(}d_{1}d_{2}\frac{(\gamma|\lambda_{2}(^{t}A.A)|)^{q}}{|\lambda_{2}(^{t}A.A)|}\bigg{)}^{\frac{1}{q}}$
which gives
$\lambda_{2}(M.^{t}M)\leq\bigg{(}\frac{d_{1}d_{2}}{|\lambda_{2}(^{t}A.A)|}\bigg{)}^{\frac{1}{q}}\gamma|\lambda_{2}(^{t}A.A)|$
By taking the limits when $q$ goes to infinity, we obtain
$\lambda_{2}(M.^{t}M)\leq\gamma|\lambda_{2}(^{t}A.A)|$
To finish the proof, we show the following:
Lemma A.3.1.
If $\lambda$ is an eigenvalue of $P$ then $\lambda^{2}$ is an eigenvalue of $M.^{t}M$.
Proof.
$P^{2}$, whose entries represents the paths of size two in the graph, is the matrix of a disconnected $d_{1}d_{2}$-regular graph. Indeed, we can remark that
[…]
and $M.^{t}M$ is symmetric. Let
$v=(v_{1},v_{2},\dots,v_{n_{1}+n_{2}})$
be an eigenvector of $P$ with eigenvalue $\lambda$. Then
$v^{\prime}=(-v_{1},-v_{2},\dots,-v_{n_{1}},v_{n_{1}+1},\dots,v_{n_{1}+n_{2}})$
is also an eigenvector with associated eigenvalue $-\lambda$. Thus, $v-v^{\prime}$ is an eigenvector of $P^{2}$ of eigenvalue $\lambda^{2}$ and this vector has $n_{2}$ zeros on the right. Because $P^{2}$ represents a disconected graph, if we reduce the dimension of this vector by $n_{2}$ (removing the zeros on the right corresponding to one connected component) we get an eigenvector of $M.^{t}M$ of eigenvalue $\lambda^{2}$. Therefore, $M.^{t}M$ has the same eigenvalues —denoted $\mu_{1}\geq\mu_{2}...\geq\mu_{n_{1}}$— as $P$, but squared. ∎
The proof works the same with $Q$ (taking $n_{1}=n_{2}$). We note $\alpha$ the bound for $|\mu_{3}(Q)|$ proven in the preceding section. $\alpha$ might refer to the bound from Theorem 1 if the graph is obtained from a bipartite regular graph of even degree or to the bound from Theorem 2 if its degree is odd. In ${}^{t}A.A$, the second largest magnitude eigenvalue is thus $|\lambda_{2}(^{t}A.A)|=(d_{1}|\mu_{3}(Q)|)^{2}\leq d_{1}^{2}\alpha^{2}$. Hence we have
$|\lambda_{2}(P)|=\sqrt{|\lambda_{2}(M.^{t}M)|}\leq\sqrt{\gamma|\lambda_{2}(A.^{t}A)|}\leq\sqrt{d_{1}d_{2}}\alpha.$
A.4 Remarks on the proofs techniques
In section A.1, we have seen that the probability for the random walk to collapse to the identity sequence is less than $(\frac{2}{d})^{m}$. If so, the probability of getting a closed walk (conditioned by the collapsing event) is then 1. When the collapse does not happen, we cannot hope for a smaller probability than $\frac{1}{n}$ to get a closed path in the graph. This is the smallest probability of a closed walk one can get in a random graph. The trace method gives then
$\mathbf{E}(|\mu_{2}|)\leq\bigg{(}n\bigg{(}\frac{1}{n}+\bigg{(}\frac{2}{d}\bigg{)}^{m}\bigg{)}-1\bigg{)}^{\frac{1}{2m}}=n^{\frac{1}{2m}}\sqrt{\frac{2}{d}}.$
With $m=\Omega(\ln n)$, we get $\mathbf{E}(|\mu_{2}|)=\mathcal{O}(d^{-\frac{1}{2}})$, which is bigger than the bound from [11] only by a constant factor. This suggests that this technique can be improved by a subtler subdivision of the probability space of $\omega$, as well as a more careful analysis of the probability of having a closed walk (specially with the condition $X_{2}^{\prime}\cap X_{4}$ for $X_{2}^{\prime}$ and $X_{4}$ defined on page A.1).
Let us observe that the term $n^{\frac{1}{2m}}$ is getting close to $1$ only when $m=\Omega(\log n)$. This is why in [6] or [3], the length of the random walk ($2m$) is logarithmic in the number of vertices. Our computations show that the optimal size of the walk should be a bit smaller; this might be because it allows us to assume that, with the overwhelming probability, at least one letter appears in $\omega$ exactly one time (the event $X_{1}$ in the proof of Theorems 1 and 3). Clearly, we cannot keep $m$ small and at the same time make the factor $n^{\frac{1}{2m}}$ close to $1$. This is an important limitation of our technique.
Our experimental results show that the second largest eigenvalue distribution measured for these graphs is much closer to that we can observe in the permutation model, at least in high dimension, and with a small field. A reasonable conjecture might be the following:
Conjecture 1.
Let $G_{k}=Sch(GL_{k}(\mathbb{F}_{q})\circlearrowleft(\mathbb{F}_{q}^{k})^{*},S)$ with $S$ a random subset of $GL_{k}(\mathbb{Z}_{q})$ and $q$ a prime number. Let $G^{\prime}_{k}$ be a $2|S|$-regular graph from the permutation model of size $q^{k}-1$. Then, as $k$ grows, the second largest eigenvalue distribution of $G$ converges to that of $G^{\prime}$.
We believe that similar statements are true for bipartite regular and biregular graphs from our construction.